OpenStudy (anonymous):

f(x)=3x-7 and g(x)=-2x-6. Find (f∘g)(4) PLEASE SHOW ALL STEPS!

OpenStudy (anonymous):

@nubeer

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@kropot72

OpenStudy (phi):

(f∘g)(4) means find g(4) by replacing x with 4 in the definition of g then put that number into f(x)

OpenStudy (anonymous):

I'm confused..

OpenStudy (phi):

can you re-write g(x)=-2x-6 by replacing x with 4 ? what do you get ?

OpenStudy (anonymous):

g(4)=-2(4)-6

OpenStudy (phi):

yes, now do the arithmetic . can you simplify -2(4)-6 ?

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

-48*

OpenStudy (phi):

-2*4 - 6 order of operations: multiply before add/subtract -2*4 is -8 now do -8 - 6= ?

OpenStudy (anonymous):

-14.

OpenStudy (phi):

when you get time, this video explains order of operations http://www.khanacademy.org/math/arithmetic/multiplication-division/order_of_operations/v/introduction-to-order-of-operations

OpenStudy (phi):

g(4)= -14 now do f( g(4) ) = f( -14)

OpenStudy (anonymous):

uhhh hahaha whoa

OpenStudy (phi):

replace x with -14 in f(x)=3x-7 simplify (multiply before add)

OpenStudy (phi):

f( g( 4)) is what (f∘g)(4) means. you find what g(4) is, and then use that number in f

OpenStudy (anonymous):

wait im confused. do i add 14 to both sides?

OpenStudy (phi):

replace x with -14 in f(x)=3x-7 same idea as for g(x): erase the x, put in a -14

OpenStudy (phi):

and 3x is short for 3 * x (3 times x)

OpenStudy (anonymous):

yes i have that. f(-14)=3(-14)-7

OpenStudy (anonymous):

do i dont do anything with the f(-14) correct? just the other stuuff?

OpenStudy (anonymous):

so -49 right? times 3*14 then subtract 7?

OpenStudy (phi):

leave f(-14) like it is. yes you get -49: 3* -14 then -7

OpenStudy (anonymous):

okay :) so now what

OpenStudy (phi):

that is the answer. But you have to show your work.

OpenStudy (anonymous):

so thats just it? i wrote it all down but it looks incomplete :o why! haha. i hate math

OpenStudy (phi):

I would say (f∘g)(4) = f( g( 4)) then g(x)=-2x-6 g(4)= fill in your work g(4)= -14 then f(x)= 3x-7 then replace x with g(4) (which you found is -14) f(-14)= -49

OpenStudy (anonymous):

okay cause mine doesn't look like that haha